The breaking strength (in pounds) of a certain new synthetic is normally distributed, with a mean of 135 and a variance of 4. The material is considered defective if the breaking strength is less than 131 pounds. What is the probability that a single, randomly selected piece of material will be defective?
Answer by stanbon(75887)

The breaking strength (in pounds) of a certain new synthetic is normally distributed, with a mean of 135 and a variance of 4. The material is considered defective if the breaking strength is less than 131 pounds. What is the probability that a single, randomly selected piece of material will be defective?
---------
z(131) = (131-135)/4 = -4/4 = -1
-------------------
P(x < 131) = P(z < -1) = normalcdf(-100,-1) = 0.1587
================
Cheers,
Stan H.
======================